2. Explain the effect of each of the following sources of error on the molarity of NaOH as determined in this experiment stating whether the concentration of NaOH obtained would be too high or too low.

a. The buret is contaminated with an acid solution.

b. The buret contains a large air bubble in the tip, which disappears in the course of the titration.

c. A small volume of the acid solution is spilled when you transfer it into the Erlenmeyer flask.

 

3. Using the molarity of the NaOH solution obtained from your titration, calculate the molarity of a 26.0 mL of CH3COOH solution that requires 32.50 mL of NaOH to reach the endpoint.

 


1
Expert's answer
2021-11-02T02:23:51-0400

2.

a)

Burette contaminated withn the acid will increase the molarity of acid if the acid is concentred than the the given acid and if it will be dilute than the given acid then it will decrease the molarity

b)

A large air bubble in the burrete will show a wrong measurement and will cause a decrease in the molarity of the given acid

c)

The spilled volume also show a decrease in the molarity of acid

3.

Let ‘M1’ be the molarity of NaOH obtained from the titration.

Volume of ‘M1’ molar NaOH used to reach end point, V1 = 32.5 mL

M1V1 = number of millimoles of NaOH

Let ‘M2’ be the molarity of CH COOH solution = ?

Volume of CH COOH solution taken, V2 = 26.0 mL

M2V2 = number of millimoles of CH COOH

Reaction: 1CH COOH + 1NaOH ----> 1CH COONa + 1H O

For this acid-base titration, at the end point we will have

M1V1 = M2V2

M1 x 32.5 = M2 x 26

Molarity of acetic acid solution: M2 = M1×32.526\frac{M1 \times 32.5}{26 }